OpenStudy (anonymous):

Let's let the equation y = 0.50(1.06)x represent the price of movie tickets for the years after 1950. You have filled in the ticket table for the prices for the years 0, 10, 20, 30, 40. Do these prices look reasonable?

OpenStudy (anonymous):

Calculator time y1 = 0.50(1.06)^0 = 0.50 y2 = 0.50(1.06)^10 = 0.90 y3 = 0.50(1.06)^20 = 1.60

OpenStudy (turingtest):

does that look like your chart meow?

OpenStudy (anonymous):

"Reasonable" is a bit grey?

OpenStudy (turingtest):

meverett has given the correct chart is it the same as the one in the question?

OpenStudy (anonymous):

Plug 0 into the equation, what do you get?

OpenStudy (anonymous):

There is no chart in the question..

OpenStudy (anonymous):

Maybe the y1 and y2 was confusing??

OpenStudy (turingtest):

sorry the "ticket table"

OpenStudy (anonymous):

Meow, Plug 0 into the equation, what do you get?

OpenStudy (anonymous):

There's no chart or table. This is all it that I posted.

OpenStudy (anonymous):

Meow, forget the chart, plug 0 into the equation, what do you get?

OpenStudy (turingtest):

did you read the question? it says " You have filled in the ticket table for the prices for the years 0, 10, 20, 30, 40. Do these prices look reasonable?" so there above is the table given in the problem

OpenStudy (turingtest):

Nevermind, just do what meverett said...

OpenStudy (anonymous):

The answer I got is 1.

OpenStudy (turingtest):

y=0.50(1.06)^(0) is not 1...

OpenStudy (anonymous):

??????????? Well that's what I typed in the calculator.... y=0.50 (1.06) ^ 0 = 1. That is how I did it and that is what I got. So I have no clue then.

OpenStudy (turingtest):

you typed it wrong I assure you what is anything to the power of 0 ???

OpenStudy (anonymous):

How could I have typed it wrong when I typed it EXACTLY how it looks into my calculator? You type it in a calculator then and tell me what you get. I'm not saying I'm right but what answer am I supposed to get when it equals 1?

OpenStudy (anonymous):

Type y = 0.50 * 1.06^0.

OpenStudy (turingtest):

I have but this doesn't even require a calculator

OpenStudy (anonymous):

There 0.5. Thank you @fewscrewsmissing.

OpenStudy (turingtest):

and what does this tell us about your "ticket table" is it reasonable?

OpenStudy (anonymous):

I don't think so.

OpenStudy (turingtest):

You are correct congrads meow, please figure out why you got the wrong number from your calculator.

OpenStudy (anonymous):

Its because I put parenthesis around them I think. But thank you too TuringTest.

OpenStudy (turingtest):

welcome